# UniVoice: Unifying Autoregressive ASR and Flow-Matching based TTS with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated promising performance in both automatic speech recognition (ASR) and text-to-speech (TTS) systems, gradually becoming the mainstream approach. However, most current approaches address these tasks separately rather than through a unified framework. This work aims to integrate these two tasks into one unified model. Although discrete speech tokenization enables joint modeling, its inherent information loss limits performance in both recognition and generation. In this work, we present UniVoice, a unified LLM framework through continuous representations that seamlessly integrates speech recognition and synthesis within a single model. Our approach combines the strengths of autoregressive modeling for speech recognition with flow matching for high-quality generation. To mitigate the inherent divergence between autoregressive and flow-matching models, we further design a dual attention mechanism, which switches between a causal mask for recognition and a bidirectional attention mask for synthesis. Furthermore, the proposed text-prefix-conditioned speech infilling method enables high-fidelity zero-shot voice cloning. Experimental results demonstrate that our method can achieve or exceed current single-task modeling methods in both ASR and zero-shot TTS tasks. This work explores new possibilities for end-to-end speech understanding and generation.
@@ Introduction
Recent years have witnessed remarkable progress in large language models (LLMs), with systems such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023), and Qwen (Bai et al. 2023) demonstrating unprecedented capabilities in understanding, generation, and reasoning. Modern LLMs have achieved broad knowledge representation and strong generalization in diverse domains.
This success has naturally been extended to speech processing, where researchers have adapted LLM frameworks to handle continuous speech signals through various discretization strategies (Chen et al. 2024a; Wang et al. 2024a; Du et al. 2023; Ma et al. 2024). The prevailing approach involves converting raw waveforms into discrete tokens using self-supervised learning representations (Hsu et al. 2021) or quantization based on neural codecs (Zeghidour et al. 2021). These tokenized representations enable the direct application of existing LLM architectures to speech tasks: for ASR tasks, Codec-ASR (Dhawan et al. 2024) performs comprehensive analysis on building ASR systems with discrete codes; for TTS tasks, systems such as VALL-E (Wang et al. 2023) and AudioLM (Borsos et al. 2023) formulate speech generation as discrete token sequence modeling. Concurrently, diffusion models (Ho, Jain, and Abbeel 2020) and flow matching models (Lipman et al. 2022; Liu, Gong, and Liu 2022) have rapidly gained prominence in speech synthesis, establishing new benchmarks for generation quality. Pioneering works like Grad-TTS (Popov et al. 2021) and F5-TTS (Chen et al. 2024c) demonstrate their exceptional ability to directly model continuous speech representations, achieving unprecedented fidelity that has made them increasingly prevalent in recent research.
Several pioneering efforts have attempted to unify ASR and TTS within single LLM frameworks using discrete representations (Wang et al. 2024a; Du et al. 2023; Tian et al. 2025). Viola (Wang et al. 2024a) converts speech utterances to discrete tokens using an offline neural codec encoder and treats all tasks as token-based sequence prediction problems, while LauraGPT (Du et al. 2023) employs a novel data representation combining continuous and discrete features for audio signals. However, such methods face inherent limitations; quantization unavoidably eliminates perceptually critical acoustic details, driving our investigation of continuous representation alternatives.
@@
We evaluated the ASR and TTS performance of our method on the LibriHeavy dataset (Kang et al. 2024). Experiments demonstrate that our approach achieves speech synthesis performance comparable to that of current state-of-the-art methods while maintaining competitive speech recognition capability.
@@
The emergence of ChatGPT has demonstrated the remarkable capabilities of LLMs, inspiring significant advancements in audio and speech processing. Recent research has successfully adapted LLMs to model discrete speech tokens, enabling high-quality speech generation and editing. Several notable approaches have emerged in this field (Wang et al. 2023; Anastassiou et al. 2024; Meng et al. 2024; Ye et al. 2025). Wang et al. (2023) introduced VALL-E, framing TTS as a conditional audio codec language modeling task with in-context learning capabilities. This work was subsequently enhanced by Chen et al. (2024a) through repetition-aware sampling and grouped codec modeling in VALL-E 2. Copet et al. (2024) developed MusicGen, utilizing delay patterns to model multiple parallel streams of audio tokens. Similarly, Peng et al. (2024) presented VoiceCraft, which combines delay patterns with causal masking for zero-shot speech editing and synthesis. Jiang et al. (2023) proposed Mega-TTS, a zero-shot TTS system that decomposes mel-spectrograms into content, timbre, prosody, and phase attributes, modeling each component according to its intrinsic properties. Yang et al. (2023) developed UniAudio, an audio foundation model capable of handling multiple generation tasks through multi-scale transformer-based codec modeling (Yu et al. 2023). Seed-TTS (Anastassiou et al. 2024) proposes utilizing both speech tokenizer based language model and diffusion acoustic modeling the first time for natural and high-quality speech generation. Similarly, CosyVoice (Du et al. 2024) proposes supervised semantic tokens for audio codec modeling and flow matching for acoustic details modeling. MaskGCT (Wang et al. 2024b) designs a masked generative codec transformer for zero-shot speech synthesis.
In addition to using LLM for modeling speech synthesis, there are also some works applying LLM for speech recognition or understanding tasks (Bai et al. 2024; Ma et al. 2024). SALMONN  (Tang et al. 2023) uses a dual encoder from Whisper speech encoder (Radford et al. 2023) and BEATS audio encoder (Chen et al. 2022b) for speech and audio understanding. SeedASR (Bai et al. 2024) is developed under the audio conditional LLM framework, utilizing the capabilities of LLM by submitting continuous speech representations, instructions, and contextual information into LLM.
@@
The diffusion (Ho, Jain, and Abbeel 2020; Song et al. 2020) models first achieved great success in the field of image generation, and then there are many works in the field of speech and audio using diffusion or flow matching (Lipman et al. 2022) to model speech and audio generation. Due to the advantages of diffusion and flow matching in modeling continuous features, diffusion- or flow-based audio generative models can often generate high-quality audio (Liu et al. 2022; Guan et al. 2024a; Anastassiou et al. 2024; Jiang et al. 2025; Wang et al. 2025b, a). Voicebox (Le et al. 2024) proposes a text-guided speech infilling task to generate masked speech segments based on surrounding audio and a provided text transcript. Matcha-TTS (Mehta et al. 2024) and VoiceFlow (Guo et al. 2024b) utilize the optimal transport conditional flow matching based on Grad-TTS (Popov et al. 2021) for high-quality and faster TTS. ReFlow-TTS (Guan et al. 2024b) models the Ordinary Differential Equation (ODE) based on Rectified Flow (Liu, Gong, and Liu 2022) for high-fidelity and efficient speech synthesis. NaturalSpeech2 (Shen et al. 2023) models zero shot speech synthesis as a conditional latent diffusion model with codec-latent embeddings and attention-based in-context modeling. FlashSpeech (Ye et al. 2024) develops a large-scale zero-shot speech synthesis with a latent consistency model and adversarial consistency training approach for efficient zero-shot speech synthesis. F5-TTS (Chen et al. 2024c) is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer for zero-shot TTS. Some works also use latent diffusion or a latent flow matching model for text-to-audio generation (Liu et al. 2023; Ghosal et al. 2023; Guan et al. 2024c).
@@
With the development of language models, they can handle various downstream tasks as a unified model in the field of text processing (Raffel et al. 2020; Brown et al. 2020). Some works have begun to explore the use of language models as the backbone of the unified model for processing speech and text tasks (Ao et al. 2021; Rubenstein et al. 2023; Zhang et al. 2023b), thus completing various related tasks in the speech modality. SpeechNet (Chen et al. 2021) and SpeechT5 (Ao et al. 2021) perform various speech tasks with an encoder-decoder model, specifically SpeechT5 needs to pre-train first and then fine-tune on subsequent tasks. Viola (Wang et al. 2024a) follows the VALL-E paradigm and integrates speech recognition, machine translation, and speech synthesis into a unified codec language model. LauraGPT (Du et al. 2023) encodes the input audio into continuous representations using an audio encoder and generates the output audio from discrete codec codes. OpusLM (Tian et al. 2025) is designed to accept and generate multistream discrete tokens in both text and speech modalities with pre-trained LLM. We propose UniVoice, which uses continuous speech representations as input and output features and combines autoregression and flow matching in one transformer, to make it more suitable for speech recognition and speech synthesis, respectively.
Note that the primary focus of this work is developing a unified speech processing model, and we do not explore speech dialogue applications (Zhang et al. 2023a; Xie and Wu 2024; Fu et al. 2024; Chen et al. 2024b) that require joint speech understanding and generation. Therefore, such tasks are beyond the scope of this paper.
@@
The objective of Flow Matching (FM)(Lipman et al. 2022) is to construct a flow that transforms a sample $X_{0}\sim p$, drawn from a source distribution $p$, into a target sample $X_{1}=\psi_{1}(X_{0})$ such that $X_{1}\sim q$ conforms to a desired distribution $q$.
@@
The ASR component processes input speech through an audio encoder module, which consists of: 1) a Whisper encoder (Radford et al. 2023) for the extraction of speech features and 2) an adapter network for semantic alignment. The inherent structure of text as a linear sequence aligns well with the standard 1D positional embeddings of the LLM, which are sufficient for text modeling and speech understanding tasks.
@@
Compared to F5-TTS (Chen et al. 2024c), we replace the adaLN-zero with the in-context modeling approach of DiT to maintain the original LLM structure. Additionally, we propose a strategic sequence organization (text prefix) to preserve the in-context learning abilities inherent in LLMs.
As illustrated in Figure 2, we introduce two variants of the TTS model for systematic comparison: (1) our primary UniVoice-TTS-infilling model that implements the voice cloning task of our UniVoice framework and (2) UniVoice-TTS-speaker, a simplified architecture designed for baseline evaluation. For UniVoice-TTS-speaker, we adopt a direct mel-spectrogram generation approach, bypassing masked infilling paradigms entirely. The system produces mel-spectrograms conditioned on input text embeddings while utilizing speaker embeddings to control vocal characteristics, enabling effective voice cloning. In this setting, for the speaker encoder, we utilize the first layer of the XLSR-53 model (Conneau et al. 2020) to extract a global embedding that effectively captures the speaker’s timbre characteristics.
@@
Following F5-TTS (Chen et al. 2024c), we also apply Classifier-Free Guidance (CFG) by dropping the condition at a certain rate during training.
@@
Following the F5-TTS approach (Chen et al. 2024c), we compute the duration ratio as:
@@
For TTS tasks, we use a bidirectional attention mask and utilize the characteristics of LLM text modeling to model the text content in TTS tasks, similar to in-context learning in VALL-E (Wang et al. 2023).
Table 1: Performance Comparisons of UniVoice and prior works. A-WER-clean, A-WER-other represents the ASR WER evaluation result on LibriSpeech test-clean dataset and LibriSpeech test-other dataset, respectively. represents that the results are reported in NaturalSpeech3(Ju et al. 2024).
@@
We conduct the training of the proposed UniVoice  using the LibriHeavy (Kang et al. 2024) dataset for both ASR and TTS training, which contains a total of about 50K hours of duration, which are sampled at 16kHz. In practice, we upsample the sample rate to 22.05kHz. For the zero-shot TTS evaluation, we use the LibriSpeech-PC test set the same as in F5-TTS (Chen et al. 2024c). For the ASR task, we directly use the LibriSpeech test-clean and test-other subsets for evaluation. We extract the 80-bin mel-spectrogram with a frame size of 1024 and a hop size of 256.
@@
We use SmolLM2-360M (Allal et al. 2025) as the underlying language model for initialization. We use the Whisper-large-v3-turbo encoder as the audio encoder, and we utilize an adaptive average pooling layer as the adapter for downsampling. For the vocoder, we use BigvGAN (Lee et al. 2022) in practice.
@@
For zero-shot TTS, we conduct a comprehensive evaluation, encompassing both objective and subjective measures, to assess the sample quality (UTMOS, CMOS), speaker similarity (SIM, SMOS) and robustness (WER). In specific, 1) for speech quality, we employ UTMOS [^1] https://github.com/sarulab-speech/UTMOS22 (Saeki et al. 2022), which is a surrogate objective metric of MOS; and we employ comparative mean option score (CMOS) to evaluate the sample naturalness subjectively; 2) for speaker similarity, we use a WavLM-large-based speaker verification model (Chen et al. 2022a) to extract speaker embeddings for calculating the cosine similarity of synthesized and ground truth speeches as SIM, and we employ similarity mean option score (SMOS) to evaluate the similarity subjectively; 3) for Word Error Rate (WER), we use an ASR model to transcribe generated speech. We use Whisper-large-v3 to compute the WER.
@@
For unified models, we compare our UniVoice  with baselines: SpeechT5 (Ao et al. 2021), LauraGPT (Du et al. 2023), OpusLM (Tian et al. 2025).
For the zero-shot TTS task, we compare our UniVoice with baselines: VALL-E (Wang et al. 2023), NaturalSpeech2 (Shen et al. 2023), F5-TTS (Chen et al. 2024c), CosyVoice (Du et al. 2024), FireRedTTS (Guo et al. 2024a), MaskGCT (Wang et al. 2024b), UniVoice-TTS. UniVoice-TTS is the model trained only for TTS using our model framework.
For the ASR task, we compare our UniVoice  with baselines: Whisper (Radford et al. 2023) series, Paraformer (Gao et al. 2022), Zipformer (Yao et al. 2023), UniVoice-ASR. UniVoice-ASR is the model trained only for ASR using our model framework.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated promising performance in both automatic speech recognition (ASR) and text-to-speech (TTS) systems, gradually becoming the mainstream approach. However, most current approaches address these tasks separately rather than through a unified framework. This work aims to integrate these two tasks into one unified model. Although discrete speech tokenization enables joint modeling, its inherent information loss limits performance in both recognition and generation. In this work, we present UniVoice, a unified LLM framework through continuous representations that seamlessly integrates speech recognition and synthesis within a single model. Our approach combines the strengths of autoregressive modeling for speech recognition with flow matching for high-quality generation. To mitigate the inherent divergence between autoregressive and flow-matching models, we further design a dual attention mechanism, which switches between a causal mask for recognition and a bidirectional attention mask for synthesis. Furthermore, the proposed text-prefix-conditioned speech infilling method enables high-fidelity zero-shot voice cloning. Experimental results demonstrate that our method can achieve or exceed current single-task modeling methods in both ASR and zero-shot TTS tasks. This work explores new possibilities for end-to-end speech understanding and generation. Code is available at https://github.com/gwh22/UniVoice.
@@ Introduction
Recent years have witnessed remarkable progress in large language models (LLMs), with systems such as GPT-4 (achiam2023gpt), LLaMA (touvron2023llama), and Qwen (bai2023qwen) demonstrating unprecedented capabilities in understanding, generation, and reasoning. Modern LLMs have achieved broad knowledge representation and strong generalization in diverse domains.
This success has naturally been extended to speech processing, where researchers have adapted LLM frameworks to handle continuous speech signals through various discretization strategies (chen2024vall; wang2024viola; du2023lauragpt; ma2024embarrassingly). The prevailing approach involves converting raw waveforms into discrete tokens using self-supervised learning representations (hsu2021hubert) or quantization based on neural codecs (zeghidour2021soundstream). These tokenized representations enable the direct application of existing LLM architectures to speech tasks: for ASR tasks, Codec-ASR (dhawan2024codec) performs comprehensive analysis on building ASR systems with discrete codes; for TTS tasks, systems such as VALL-E (wang2023neural) and AudioLM (borsos2023audiolm) formulate speech generation as discrete token sequence modeling. Concurrently, diffusion models (ho2020denoising) and flow matching models (lipman2022flow; liu2022flow) have rapidly gained prominence in speech synthesis, establishing new benchmarks for generation quality. Pioneering works like Grad-TTS (popov2021grad) and F5-TTS (chen2024f5) demonstrate their exceptional ability to directly model continuous speech representations, achieving unprecedented fidelity that has made them increasingly prevalent in recent research.
Several pioneering efforts have attempted to unify ASR and TTS within single LLM frameworks using discrete representations (wang2024viola; du2023lauragpt; tian2025opuslm). Viola (wang2024viola) converts speech utterances to discrete tokens using an offline neural codec encoder and treats all tasks as token-based sequence prediction problems, while LauraGPT (du2023lauragpt) employs a novel data representation combining continuous and discrete features for audio signals. However, such methods face inherent limitations; quantization unavoidably eliminates perceptually critical acoustic details, driving our investigation of continuous representation alternatives.
@@
We evaluated the ASR and TTS performance of our method on the LibriHeavy dataset (kang2024libriheavy). Experiments demonstrate that our approach achieves speech synthesis performance comparable to that of current state-of-the-art methods while maintaining competitive speech recognition capability.
@@
The emergence of ChatGPT has demonstrated the remarkable capabilities of LLMs, inspiring significant advancements in audio and speech processing. Recent research has successfully adapted LLMs to model discrete speech tokens, enabling high-quality speech generation and editing. Several notable approaches have emerged in this field (wang2023neural; anastassiou2024seed; meng2024melle; ye2025llasa). wang2023neural introduced VALL-E, framing TTS as a conditional audio codec language modeling task with in-context learning capabilities. This work was subsequently enhanced by chen2024vall through repetition-aware sampling and grouped codec modeling in VALL-E 2. copet2024simple developed MusicGen, utilizing delay patterns to model multiple parallel streams of audio tokens. Similarly, peng2024voicecraft presented VoiceCraft, which combines delay patterns with causal masking for zero-shot speech editing and synthesis. jiang2023mega proposed Mega-TTS, a zero-shot TTS system that decomposes mel-spectrograms into content, timbre, prosody, and phase attributes, modeling each component according to its intrinsic properties. yang2023uniaudio developed UniAudio, an audio foundation model capable of handling multiple generation tasks through multi-scale transformer-based codec modeling (yu2023megabyte). Seed-TTS (anastassiou2024seed) proposes utilizing both speech tokenizer based language model and diffusion acoustic modeling the first time for natural and high-quality speech generation. Similarly, CosyVoice (du2024cosyvoice) proposes supervised semantic tokens for audio codec modeling and flow matching for acoustic details modeling. MaskGCT (wang2024maskgct) designs a masked generative codec transformer for zero-shot speech synthesis.
In addition to using LLM for modeling speech synthesis, there are also some works applying LLM for speech recognition or understanding tasks (bai2024seed; ma2024embarrassingly). SALMONN  (tang2023salmonn) uses a dual encoder from Whisper speech encoder (radford2023robust) and BEATS audio encoder (chen2022beats) for speech and audio understanding. SeedASR (bai2024seed) is developed under the audio conditional LLM framework, utilizing the capabilities of LLM by submitting continuous speech representations, instructions, and contextual information into LLM.
@@
The diffusion (ho2020denoising; song2020score) models first achieved great success in the field of image generation, and then there are many works in the field of speech and audio using diffusion or flow matching (lipman2022flow) to model speech and audio generation. Due to the advantages of diffusion and flow matching in modeling continuous features, diffusion- or flow-based audio generative models can often generate high-quality audio (liu2022diffsinger; guan2024mmtts; anastassiou2024seed; jiang2025sparse; wang2025spark; wang2025discl). Voicebox (le2024voicebox) proposes a text-guided speech infilling task to generate masked speech segments based on surrounding audio and a provided text transcript. Matcha-TTS (mehta2024matcha) and VoiceFlow (guo2024voiceflow) utilize the optimal transport conditional flow matching based on Grad-TTS (popov2021grad) for high-quality and faster TTS. ReFlow-TTS (guan2024reflow) models the Ordinary Differential Equation (ODE) based on Rectified Flow (liu2022flow) for high-fidelity and efficient speech synthesis. NaturalSpeech2 (shen2023naturalspeech) models zero shot speech synthesis as a conditional latent diffusion model with codec-latent embeddings and attention-based in-context modeling. FlashSpeech (ye2024flashspeech) develops a large-scale zero-shot speech synthesis with a latent consistency model and adversarial consistency training approach for efficient zero-shot speech synthesis. F5-TTS (chen2024f5) is a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer for zero-shot TTS. Some works also use latent diffusion or a latent flow matching model for text-to-audio generation (liu2023audioldm; ghosal2023text; guan24b_lafma).
@@
With the development of language models, they can handle various downstream tasks as a unified model in the field of text processing (raffel2020t5; brown2020gpt3). Some works have begun to explore the use of language models as the backbone of the unified model for processing speech and text tasks (ao2021speecht5; rubenstein2023audiopalm; zhang2023google), thus completing various related tasks in the speech modality. SpeechNet (chen2021speechnet) and SpeechT5 (ao2021speecht5) perform various speech tasks with an encoder-decoder model, specifically SpeechT5 needs to pre-train first and then fine-tune on subsequent tasks. Viola (wang2024viola) follows the VALL-E paradigm and integrates speech recognition, machine translation, and speech synthesis into a unified codec language model. LauraGPT (du2023lauragpt) encodes the input audio into continuous representations using an audio encoder and generates the output audio from discrete codec codes. OpusLM (tian2025opuslm) is designed to accept and generate multistream discrete tokens in both text and speech modalities with pre-trained LLM. We propose UniVoice, which uses continuous speech representations as input and output features and combines autoregression and flow matching in one transformer, to make it more suitable for speech recognition and speech synthesis, respectively.
Note that the primary focus of this work is developing a unified speech processing model, and we do not explore speech dialogue applications (zhang2023speechgpt; xie2024mini; fu2024vita; chen2024slam) that require joint speech understanding and generation. Therefore, such tasks are beyond the scope of this paper.
@@
The objective of Flow Matching (FM)(lipman2022flow) is to construct a flow that transforms a sample $X_{0}\sim p$, drawn from a source distribution $p$, into a target sample $X_{1}=\psi_{1}(X_{0})$ such that $X_{1}\sim q$ conforms to a desired distribution $q$.
@@
The ASR component processes input speech through an audio encoder module, which consists of: 1) a Whisper encoder (radford2023robust) for the extraction of speech features and 2) an adapter network for semantic alignment. The inherent structure of text as a linear sequence aligns well with the standard 1D positional embeddings of the LLM, which are sufficient for text modeling and speech understanding tasks.
@@
Compared to F5-TTS (chen2024f5), we replace the adaLN-zero with the in-context modeling approach of DiT to maintain the original LLM structure. Additionally, we propose a strategic sequence organization (text prefix) to preserve the in-context learning abilities inherent in LLMs.
As illustrated in Figure 2, we introduce two variants of the TTS model for systematic comparison: (1) our primary UniVoice-TTS-infilling model that implements the voice cloning task of our UniVoice framework and (2) UniVoice-TTS-speaker, a simplified architecture designed for baseline evaluation. For UniVoice-TTS-speaker, we adopt a direct mel-spectrogram generation approach, bypassing masked infilling paradigms entirely. The system produces mel-spectrograms conditioned on input text embeddings while utilizing speaker embeddings to control vocal characteristics, enabling effective voice cloning. In this setting, for the speaker encoder, we utilize the first layer of the XLSR-53 model (conneau2020unsupervised) to extract a global embedding that effectively captures the speaker’s timbre characteristics.
@@
Following F5-TTS (chen2024f5), we also apply Classifier-Free Guidance (CFG) by dropping the condition at a certain rate during training.
@@
Following the F5-TTS approach (chen2024f5), we compute the duration ratio as:
@@
For TTS tasks, we use a bidirectional attention mask and utilize the characteristics of LLM text modeling to model the text content in TTS tasks, similar to in-context learning in VALL-E (wang2023neural).
Table 1: Performance Comparisons of UniVoice and prior works. A-WER-clean, A-WER-other represents the ASR WER evaluation result on LibriSpeech test-clean dataset and LibriSpeech test-other dataset, respectively. represents that the results are reported in NaturalSpeech3(ju2024naturalspeech).
@@
We conduct the training of the proposed UniVoice  using the LibriHeavy (kang2024libriheavy) dataset for both ASR and TTS training, which contains a total of about 50K hours of duration, which are sampled at 16kHz. In practice, we upsample the sample rate to 22.05kHz. For the zero-shot TTS evaluation, we use the LibriSpeech-PC test set the same as in F5-TTS (chen2024f5). For the ASR task, we directly use the LibriSpeech test-clean and test-other subsets for evaluation. We extract the 80-bin mel-spectrogram with a frame size of 1024 and a hop size of 256.
@@
We use SmolLM2-360M (allal2025smollm2smolgoesbig) as the underlying language model for initialization. We use the Whisper-large-v3-turbo encoder as the audio encoder, and we utilize an adaptive average pooling layer as the adapter for downsampling. For the vocoder, we use BigvGAN (lee2022bigvgan) in practice.
@@
For zero-shot TTS, we conduct a comprehensive evaluation, encompassing both objective and subjective measures, to assess the sample quality (UTMOS, CMOS), speaker similarity (SIM, SMOS) and robustness (WER). In specific, 1) for speech quality, we employ UTMOS [^1] https://github.com/sarulab-speech/UTMOS22 (saeki2022utmos), which is a surrogate objective metric of MOS; and we employ comparative mean option score (CMOS) to evaluate the sample naturalness subjectively; 2) for speaker similarity, we use a WavLM-large-based speaker verification model (chen2022wavlm) to extract speaker embeddings for calculating the cosine similarity of synthesized and ground truth speeches as SIM, and we employ similarity mean option score (SMOS) to evaluate the similarity subjectively; 3) for Word Error Rate (WER), we use an ASR model to transcribe generated speech. We use Whisper-large-v3 to compute the WER.
@@
For unified models, we compare our UniVoice  with baselines: SpeechT5 (ao2021speecht5), LauraGPT (du2023lauragpt), OpusLM (tian2025opuslm).
For the zero-shot TTS task, we compare our UniVoice with baselines: VALL-E (wang2023neural), NaturalSpeech2 (shen2023naturalspeech), F5-TTS (chen2024f5), CosyVoice (du2024cosyvoice), FireRedTTS (guo2024fireredtts), MaskGCT (wang2024maskgct), UniVoice-TTS. UniVoice-TTS is the model trained only for TTS using our model framework.
For the ASR task, we compare our UniVoice  with baselines: Whisper (radford2023robust) series, Paraformer (gao2022paraformer), Zipformer (yao2023zipformer), UniVoice-ASR. UniVoice-ASR is the model trained only for ASR using our model framework.